OpenStudy (anonymous):

f(x)=-5x^2+7x+4 I need the x-coordinate of the vertex of the parabola defined by this function

OpenStudy (anonymous):

what class is this for?

OpenStudy (anonymous):

you can take the derivative and set it to 0 and solve for x

OpenStudy (anonymous):

If you haven't learned derivatives, it is in your best interest to convert this to vertex form. You are currently in\[ax^2+bx+c\] but you want it in \[a(x-h)^2+k\] where \[(h,k)\] is your vertex

OpenStudy (anonymous):

Welcome to openstudy by the way

Directrix (directrix):

The x-coordinate of the vertex of a parabola is x = -b/(2*a). In f(x) = f(x)=-5x^2+7x+4, the a comes from the coefficient of the x^2 term and is- 5. The b is the coefficient of the x term and is 7. x = -b/(2*a) = -7/(2 * -5) = ? @klapratt72